A play yard is 19ft by 32 ft. A fence is to be built around the yard. How many feet of fencing will be needed? If fencing costs $6.60 per foot, what will be the cost of the fence?

Answer by mananth(16949)

Perimeter = 2(L+W)
L =32
w=19
perimeter = 2*(32+19)
perimeter = 102
cost is $6.60 /ft
total cost = 102 * 6.60
=$673.20